Solve the equation for x:
x^2+4x+4=7
I am not sure where to start with this one. It is strange because of the extra number. How do I do this problem? PLEASE HELP! THANK YOU!
ac

Answer by checkley71(8403)

X^2+4X+4=7
X^2+4X-3=0 USING THE QUADRATIC EQUATION X=(-B+-SQRT[B^2-4AC])/2A
WE GET
X=(-4+-SQRT[4^2-4*1(-3])/2*1
X=(-4+-SQRT[16+12])/2
X=(-4+-SQRT[28])/2
X=(-4+-5.2915)/2
X=(-4+5.2915)/2
X=(1.2915/2
X=.65575 SOLUTION
X=(-4-5.2915)/2
X=-9.2915/2
X=-4.64575 SOLUTION